OpenStudy (anonymous):

In comparing the graphs of y = 3^x and y = 2^x, which statement is TRUE? Select one: a. The graph of y = 3^x is further away form the y-axis than y = 2^x. b. The graph of y = 2^x is further away form the y-axis than y = 3^x. c. The graph of y = 2^x increases at a slower rate than y = 3^x. d. The graph of y = 3^x increases at a slower rate than y = 2^x.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

B. When you graph it, y=2^x is farther from y=3^x